# Adaptive Alpha Weighting with PPO: Enhancing Prompt-Based LLM-Generated Alphas in Quant Trading

### v1

@@ 5.1 Impact of Different Alpha Selection Settings on Trading Performance
Table 9 presents the results when using low-correlation alphas. Comparing the PPO strategy using all 50 alphas with the version that applies low-correlation filtering, we observe that reducing redundancy among alphas affects stocks differently. For Apple and Tencent, performance declines slightly in terms of returns or risk-adjusted metrics, suggesting that some removed alphas may have contributed useful information. In contrast, HSBC, Pepsi, and Toyota show modest improvements, indicating that the elimination of highly correlated alphas can slightly enhance the signal quality for these stocks. In general, the low-correlation approach maintains the ability of the strategy to outperform market benchmarks while slightly increasing variability for certain stocks, reflecting a trade-off between signal diversity and stability.
Table 10 shows the performance using high-contribution alphas. Specifically, Apple, Pepsi, and Tencent experience slight declines in cumulative returns and Sharpe ratios, along with higher drawdowns, suggesting that focusing solely on the top-contributing alphas may reduce the diversity of useful signals for these stocks. In contrast, HSBC and Toyota see improvements in both return and risk-adjusted performance, indicating that high-contribution alphas provide particularly strong signals for these assets. Overall, while the high-contribution approach maintains the ability to outperform market benchmarks, it does not consistently improve performance across all stocks and may sacrifice stability for certain assets.
Table 11 presents the performance of the PPO-adjusted strategy when using thirty randomly selected alphas. Apple continues to exhibit strong performance, achieving the highest cumulative return and Sharpe ratio among all five stocks. In contrast, Toyota still performs the worst in this setting. Notably, HSBC shows improved performance compared to the scenario using all fifty alphas in terms of return and Sharpe ratio. Overall, despite the reduced number of alphas, all stocks still outperform their respective market benchmarks.
@@ 5.2 Impact of Prompt Information on Trading Performance
For simplicity, the analysis focuses solely on Apple. In the partial information scenario, only data from Apple is provided. Table 13 demonstrates the PPO strategy performance under different scenarios. Based on the results, we observe that across all scenarios, formulaic alphas generated with varying prompt information produce strong performance under the proposed RL framework, with cumulative returns exceeding the market benchmark and Sharpe ratios above one, demonstrating the robustness and effectiveness of the framework.
@@ 5.3 Impact of Sentiment on Trading Performance
Table 14 to 18 report the performance of the proposed PPO strategy for all five companies under different sentiment settings. The results suggest that incorporating sentiment features within the PPO framework does not lead to substantial differences in trading performance. Across all three sentiment settings for each target company, the proposed strategy consistently delivers strong results.
Table 14: Performance of PPO-Adjusted Strategy on Apple under Different Sentiment Settings. Mean (Standard Deviation).
Metric Sentiment of Target Company Only Sentiments of All Companies No Sentiment
Cumulative Return 1.2499 (0.8023) 1.4797 ( 0.1156) 1.4459 (0.1098)
@@
Drawdown -0.0133 (0.0013) -0.0084 (0.0019) -0.0110 (0.0020)
Market Return 0.3476 (S&P 500) 0.3476 (S&P 500) 0.3476 (S&P 500)
Table 15: Performance of PPO-Adjusted Strategy on HSBC under Different Sentiment Settings. Mean (Standard Deviation).
Metric Sentiment of Target Company Only Sentiments of All Companies No Sentiment
Cumulative Return 0.6093 (0.0455) 0.4460 (0.0525) 0.5955 (0.0481)
@@
Drawdown -0.1014 (0.0239) -0.1597 (0.0311) -0.1489 (0.0283)
Market Return 0.0033 (Hang Seng) 0.0033 (Hang Seng) 0.0033 (Hang Seng)
Table 16: Performance of PPO-Adjusted Strategy on Pepsi under Different Sentiment Settings. Mean (Standard Deviation).
Metric Sentiment of Target Company Only Sentiments of All Companies No Sentiment
Cumulative Return 0.4124 (0.0396) 0.7643 (0.0514) 0.7644 (0.0437)
@@
Drawdown -0.0454 (0.0111) -0.0051 (0.0070) -0.0047 (0.0064)
Market Return 0.3476 (S&P 500) 0.3476 (S&P 500) 0.3476 (S&P 500)
Table 17: Performance of PPO-Adjusted Strategy on Tencent under Different Sentiment Settings. Mean (Standard Deviation).
Metric Sentiment of Target Company Only Sentiments of All Companies No Sentiment
Cumulative Return 0.6136 (0.0486) 0.7163 (0.0436) 0.7067 (0.0513)
@@
Drawdown -0.0515 (0.0120) -0.0504 (0.0132) -0.0484 (0.0112)
Market Return 0.0033 (Hang Seng) 0.0033 (Hang Seng) 0.0033 (Hang Seng)
Table 18: Performance of PPO-Adjusted Strategy on Toyota under Different Sentiment Settings. Mean (Standard Deviation).
Metric Sentiment of Target Company Only Sentiments of All Companies No Sentiment
Cumulative Return 0.0272 (0.0554) 0.1088 (0.0622) 0.0102 (0.0512)
@@
Drawdown -0.2735 (0.0413) -0.2203 (0.0370) -0.2651 (0.0395)
Market Return 0.4081 (Nikkei 225) 0.4081 (Nikkei 225) 0.4081 (Nikkei 225)
@@ 6 Conclusion
This research explores the use of LLMs for generating formulaic alphas and applies PPO to optimize their weights for trading decisions. Experimental results show that the PPO-adjusted strategy achieves significantly higher cumulative returns than traditional benchmarks such as the S&P 500, Hang Seng, and Nikkei 225 for most target stocks. In addition to higher returns, the strategy also demonstrates strong risk-adjusted performance, as reflected in consistently high Sharpe ratios across companies like Apple and Pepsi. Moreover, the strategy maintains relatively low drawdown levels, indicating better downside protection compared to market indices. Additionally, the PPO strategy outperforms an equal-weighted alpha portfolio for most of the stocks.
Reducing the number of alphas leads to slight improvements in both cumulative return and Sharpe ratio for certain stocks. However, for others, their performance may worsen after reducing the number of alphas. This indicates that selecting a well-balanced set of informative and diverse alphas can enhance both stability and profitability. The results underscore the potential of combining LLMs with reinforcement learning to create effective and interpretable trading strategies. Future work will explore alternative LLM architectures, additional reinforcement learning methods, and broader applications across different markets and asset classes.

### v2

@@ 5.4 Impact of Different Alpha Selection Settings on Trading Performance
Table 16: Comparison of PPO Strategies across different Alpha Selection Methods. The best value for each stock under each metric is bolded.
Metric / Alpha Strategy Apple HSBC Pepsi Tencent Toyota
Cumulative Return
   Low-Correlation 1.6641 (0.1755) 0.4983 (0.0660) 0.6319 (0.0775) 0.5956 (0.0722) 0.0842 (0.0579)
   High-Contribution 1.6072 (0.1093) 0.5199 (0.0991) 0.6023 (0.0441) 0.5634 (0.1200) 0.0741 (0.0395)
   30 Random 1.7276 (0.0647) 0.4853 (0.0451) 0.6176 (0.0225) 0.6041 (0.0755) 0.0410 (0.0291)
Sharpe Ratio
   Low-Correlation 2.0002 (0.0692) 0.8388 (0.0776) 1.4259 (0.0622) 1.1191 (0.1018) 0.2072 (0.1083)
   High-Contribution 1.9890 (0.0626) 0.8791 (0.1215) 1.3996 (0.0296) 1.0620 (0.1735) 0.1894 (0.0765)
   30 Random 2.0133 (0.0255) 0.8399 (0.0534) 1.4271 (0.0367) 1.1129 (0.1132) 0.1238 (0.0565)
Max Drawdown
   Low-Correlation 0.0113-0.0113 (0.0016) $\mathbf{-0.2210}$ (0.0121) $\mathbf{-0.0066}$ (0.0002) $\mathbf{-0.0810}$ (0.0183) $\mathbf{-0.2344}$ (0.0288)
   High-Contribution 0.0113-0.0113 (0.0016) 0.2286-0.2286 (0.0253) 0.0076-0.0076 (0.0022) 0.0919-0.0919 (0.0263) 0.2625-0.2625 (0.0217)
   30 Random $\mathbf{-0.0109}$ (0.0009) 0.2248-0.2248 (0.0061) 0.0068-0.0068 (0.0006) 0.0937-0.0937 (0.0307) 0.2664-0.2664 (0.0137)
Note: Values represent the mean across test runs with standard deviation in parentheses.
@@
Table 16 presents the results when using different alpha selection methods. First of all, when comparing the PPO strategy using all 50 alphas with the version that applies low-correlation filtering, we observe that reducing redundancy among alphas affects stocks differently. For Apple and Tencent, performance declines slightly in terms of returns or risk-adjusted metrics, suggesting that some removed alphas may have contributed useful information. In contrast, HSBC, Pepsi, and Toyota show modest improvements, indicating that the elimination of highly correlated alphas can slightly enhance the signal quality for these stocks. In general, the low-correlation approach maintains the ability of the strategy to outperform market benchmarks while slightly increasing variability for certain stocks, reflecting a trade-off between signal diversity and stability.
When it comes to performance using high-contribution alphas, Apple, Pepsi, and Tencent experience slight declines in cumulative returns and Sharpe ratios, along with higher drawdowns, suggesting that focusing solely on the top-contributing alphas may reduce the diversity of useful signals for these stocks. In contrast, HSBC and Toyota see improvements in both return and risk-adjusted performance, indicating that high-contribution alphas provide particularly strong signals for these assets. Overall, while the high-contribution approach maintains the ability to outperform market benchmarks, it does not consistently improve performance across all stocks and may sacrifice stability for certain assets.
Furthermore, when using thirty randomly selected alphas, Apple continues to exhibit strong performance, achieving the highest cumulative return and Sharpe ratio among all five stocks. In contrast, Toyota still performs the worst in this setting. Notably, HSBC shows improved performance compared to the scenario using all fifty alphas in terms of return and Sharpe ratio.
@@ 5.5 Impact of Prompt Information on Trading Performance
For simplicity, the analysis focuses solely on Apple. In the partial information scenario, only data from Apple is provided. Table 17 demonstrates the PPO strategy performance under different scenarios. Based on the results, we observe that across all scenarios, formulaic alphas generated with varying prompt information produce strong performance under the proposed RL framework, with Sharpe ratios above one, demonstrating the robustness and effectiveness of the framework.
@@ 5.6 Impact of Sentiment on Trading Performance
Table 18 reports the performance of the proposed PPO strategy for all five companies under different sentiment settings. The results suggest that incorporating sentiment features within the PPO framework does not lead to substantial differences in trading performance. Across all three sentiment settings for each target company, the proposed strategy consistently delivers strong results.
Table 18: Performance Comparison of PPO-Adjusted Strategy across Different Sentiment Settings. Values are Mean (Standard Deviation).
Stock Metric Target Only All Companies No Sentiment
Apple Cumulative Return 1.2499 (0.8023) 1.4797 (0.1156) 1.4459 (0.1098)
@@
Max Drawdown 0.0133-0.0133 (0.0013) 0.0084-0.0084 (0.0019) 0.0110-0.0110 (0.0020)
HSBC Cumulative Return 0.6093 (0.0455) 0.4460 (0.0525) 0.5955 (0.0481)
@@
Max Drawdown 0.1014-0.1014 (0.0239) 0.1597-0.1597 (0.0311) 0.1489-0.1489 (0.0283)
Pepsi Cumulative Return 0.4124 (0.0396) 0.7643 (0.0514) 0.7644 (0.0437)
@@
Max Drawdown 0.0454-0.0454 (0.0111) 0.0051-0.0051 (0.0070) 0.0047-0.0047 (0.0064)
Tencent Cumulative Return 0.6136 (0.0486) 0.7163 (0.0436) 0.7067 (0.0513)
@@
Max Drawdown 0.0515-0.0515 (0.0120) 0.0504-0.0504 (0.0132) 0.0484-0.0484 (0.0112)
Toyota Cumulative Return 0.0272 (0.0554) 0.1088 (0.0622) 0.0102 (0.0512)
@@
Max Drawdown 0.2735-0.2735 (0.0413) 0.2203-0.2203 (0.0370) 0.2651-0.2651 (0.0395)
@@ 6 Conclusion
This study explores the use of large language models to generate formulaic alphas and applies PPO to dynamically optimize their weights for trading decisions. Empirical results show that, while the PPO-adjusted strategy does not consistently achieve the highest cumulative returns compared with traditional benchmarks such as buy-and-hold and momentum strategies, it demonstrates clear advantages in terms of risk-adjusted performance. Specifically, the PPO strategy achieves higher Sharpe ratios across most target stocks, indicating more stable return profiles relative to volatility. In addition, the strategy consistently exhibits lower maximum drawdowns, suggesting improved downside risk control and enhanced robustness during adverse market conditions. Compared with an equal-weighted alpha portfolio, the PPO-based approach further highlights the benefit of adaptive weight allocation in balancing return and risk. Overall, these findings suggest that integrating LLM-generated alphas with reinforcement learning offers a promising framework for constructing more risk-aware and resilient trading strategies, particularly in settings where drawdown control and return stability are prioritized over absolute returns.
Furthermore, the ablation analysis further clarifies the roles of both alpha construction and reinforcement learning algorithms within the proposed framework. For most of the selected stocks, portfolios built on LLM-generated alphas consistently exhibit superior performance compared to those based on human-crafted alphas. In terms of optimization methods, alternative reinforcement learning algorithms, including SAC, TD3, and A2C, achieve performance comparable to PPO under the same environmental setup.
Besides, reducing the number of alphas leads to slight improvements in both cumulative return and Sharpe ratio for certain stocks. However, for others, their performance may worsen after reducing the number of alphas. This indicates that selecting a well-balanced set of informative and diverse alphas can enhance both stability and profitability. The results underscore the potential of combining LLMs with reinforcement learning to create effective and interpretable trading strategies.
This study has several limitations. First, the analysis is conducted on only ten stocks, which may limit the generalizability of the findings. Future work will further extend the analysis to a larger and more diverse set of stocks. Second, due to limited access to high-frequency market data, the current experiments are based on daily observations, which constrains the practical applicability of the proposed LLM-based framework in real-world trading environments where higher-frequency information and richer factor signals are often available. Future studies will consider high-frequency data and incorporate a broader range of explanatory factors, including macroeconomic variables and firm-level characteristics, to better capture market dynamics. In addition, future work will explore alternative LLM architectures, and broader applications across different markets and asset classes.